# The Forest Behind the Tree: Revealing Hidden Smart Home Communication Patterns

### v3

@@ I Introduction
In this paper, we present a fully automated framework to comprehensively uncover the communication patterns of a Smart Home device. For each type of interaction with the device, it first observes the traffic flows that occur in an unconstrained network and builds a set of flow descriptors from them. Such descriptors contain Internet and transport-layer information, as well as selected, available application-layer data, such as domain names. The system then iteratively blocks flows corresponding to certain descriptors and repeats the interaction with the intent of making new flows appear. The algorithm selecting which flows to block and the sequence of blocking experiments is not straightforward. Our aim is to discover as many hidden flows as possible but keep the number of experiments under control. To do this efficiently and not end up in an infinite loop where the same patterns appear over and over again, we store the found descriptors in a tree that we traverse using a breadth-first strategy, and prune using an ad hoc heuristic. The result is a set of multi-level, tree-shaped profiles (one for each type of interaction) for the device that describe the device’s default and alternative communication patterns that we then explore to assess the robustness of the device’s network communications.
@@
  • We develop and implement a framework, modular and easily extensible, which instruments Smart Home devices and automatically extracts multi-level profiles for them. The core of the framework is a new algorithm that, for a given device interaction, builds an event signature, i.e., the description of flows appearing for that interaction, and then iteratively blocks previously observed flows, until the interaction fails or no new flow is discovered.
@@
  • We publish the source code of our signature extraction algorithm and our experimental framework, including the captures of the testbed’s traffic. Those will be made available after the anonymous review process.
@@ II Problem statement and scope
The terms IoT and Smart Home are quite fuzzily defined in the literature. Our work focuses on domestic devices, augmented with IoT capabilities, which exhibit goal-oriented, easily identifiable functions, e.g. power plugs, light bulbs, and cameras. Devices such as Smart TVs and speaker hubs are considered out of scope, as their capabilities have broadened tremendously over the years, including the possibility to install third-party apps to add new features. This makes the network behavior of such devices very complex and customizable, rendering an automated analysis with our framework practically unfeasible.
@@ III-A General workflow
Fig. 2 gives an overview on the workflow of our methodology to obtain a profile for a Smart Home device under a certain type of interaction. An interaction can require the usage of a triggering source, such as a companion app installed on a smartphone. In practice, a Smart Home device may support different interactions (e.g., for a smart lamp: switching on/off, changing its color, etc.) and triggering sources, and the workflow described here must be repeated for each of them. The workflow consists of multiple automated steps summarized below.
@@ III-B Step 1: Traffic capturing
Similarly to PingPong [11] and IoTAthena [15], we start capturing the network traffic, perform the interaction, then stop the traffic capture after a predefined duration $d$. We repeat each interaction and the traffic capturing $m$ times. Our framework automates interactions triggered by a companion app by leveraging the Android Debug Bridge (ADB) tool [22] to generate touch events on the smartphone running the companion app. We also consider power-cycling the Smart home device as an interaction (called boot interaction in the following). To automate the boot interaction, we plug the device in a smart power outlet controlled by our framework.
@@ III-D2 Tree traversal and pruning
Secondly, we prune branches of the tree which will likely not provide new information, i.e. no new Flow IDs, by applying an ad hoc pruning heuristic. To determine which heuristic we can apply without losing any information, we ran preliminary experiments with one of our testbed’s devices, the TP-Link HS110 smart plug [27]. We observed that, for any two nodes with the same Flow ID, their children were always identical. We conclude it is likely that no new information will be found by processing a node identical to one which has already been processed, regardless of its position in the tree. Therefore, we adopt the node pruning heuristic, illustrated in Fig. 3: a node is pruned if an equivalent node, i.e., a node with the same Flow ID, has already been processed.
@@ III-D3 How to block packets matching a Flow ID
As explained, a Flow ID comprises traffic features from multiple networking layers, such as hostnames, port numbers, and application layer information. The presence of application layer information in the Flow ID requires a firewall that is able to match and reject traffic based on information found at that layer. A simple firewall, such as the default Linux kernel firewall NFTables [28] (successor of the well-known IPTables) is not sufficient. Instead, we leverage the open-source Smart Home firewall by De Keersmaeker et al. [9]. It is based on NFTables, while enhancing its capabilities to match additional application-layer protocols. However, this firewall can only work with allow rules, while our profiling algorithm is based on blocking rules. Therefore, we modified the firewall, turning it into a deny-list firewall. In our prototype implementation of the profiling algorithm, we translate the list of Flow IDs into firewall blocking rules and let the firewall enforce them for the next batch of experiments. The code of the modified firewall will be made publicly available.
@@ IV-A Experiment setup
In a real Smart Home network, a user could choose among various apps to control their devices, including the device’s official companion app (from its manufacturer or vendor), a third-party app, or a Smart Home automation platform which provides unified control over devices from different manufacturers. To cover such cases, for all interactions other than booting, we derive up to three usage scenarios:
@@
Woox R5024 (Tuya) [35] / boot 4
@@
Xiaomi MJSXJ02CM [36] / boot 5
@@
[…]
@@
Alecto Smart-Bulb10 (Tuya) [40] / boot 9
@@
[…]
The complete device references are listed in Table I, along with the corresponding smartphone apps and interactions instrumented. As explained, we only consider IP traffic, whether over Ethernet or Wi-Fi. Therefore, for the two devices communicating using Zigbee, we profile the traffic issued by their hub, i.e. the Hue bridge for the Hue lamp, and the SmartThings hub for the SmartThings Outlet, respectively. For the boot interaction experiments over those devices, we synchronously power-cycle both the device itself and its hub.
@@ IV-B Example signature tree
Figure 5: Event signature tree for the TP-Link HS110’s toggle event. Bidirectional flows are represented by a double-headed arrow, whilst a single-headed arrow is used for unidirectional flows. When indicated, the port is associated with its host.
@@
    between the phone and the plug’s TCP port $9$;
@@
    between the phone and the plug’s UDP port $9$;
@@
    from the phone to the broadcast address’ (255.255.255.255) UDP port $9$;
@@ IV-C Assessing device event robustness
The experiments in our testbed of devices result in the extraction of 254 unique Flow IDs, of which 70 are hidden (i.e. 27.56%). Their distribution over the 36 tested event configurations is shown in Fig. 11 in the appendix. Hidden Flow IDs embody the alternative communication strategies that a device might use if the default one fails, effectively enhancing the device’s robustness to network instability (remember that we only keep Flow IDs associated with successful event executions). As such, we rebrand the count of hidden Flow IDs as the robustness score, and compute it for each tested event; Fig. 6 shows the results.
Figure 6: Robustness score for all events.
@@
Figure 7: Robustness score for all device interactions, grouped per device type, app, and manufacturer. Each marker represents the robustness score of one event.
@@ IV-D A closer look at DNS data
0.4Refer to caption Refer to caption OcFFL
Figure 8: Count of unique domain names / DNS resolvers contacted per device issuing the event boot
@@ IV-D2 DNS resolvers
Out of the 17 DNS resolvers contacted, only two are hidden resolvers. One of them is simply the LAN gateway (192.168.1.1), meaning the only real new DNS resolver discovered at a level deeper than one is the Chinese public DNS resolver 114.114.114.114, contacted by the Xiaomi camera when the LAN gateway is not responsive. This result shows that, regarding domain name resolving, Smart Home devices still lack robustness. Indeed, every device should try contacting a backup resolver if the primary one fails, as this has been made easy thanks to publicly available resolvers such as Cloudflare’s 1.1.1.1 or Google’s $8.$
@@ IV-E Analysis of the node pruning heuristic
Figure 9: Count of pruned Flow IDs, per depth and per event
@@ IV-F Comparison with related work
TP-Link HS110 boot & toggle 5 4 use1-api.tplinkra.com euw1-api.tplinkra.com euw1-device-telemetry-gw.iot.i.tplinknbu n-wap.tplinkcloud.com n-devs.tplinkcloud.com n-deventry.tplinkcloud.com
SmartThings Outlet boot 4 3 / hub.smartthings.com usher.connect.smartthings.com connectivity.smartthings.com dc2-eu01-euwest1.connect.smartthings.com fw-update2.smartthings.com
Philips Hue lamp boot 5 4 / mqtt-eu-01.iot.meethue.com time.meethue.com ntp2.aliyun.com ntp4.aliyun.com time4.google.com diagnostics.meethue.com ecdinterface.philips.com
@@
A direct comparison of the domain names beyond the mere numbers is unfortunately not possible. This is because Mandalari et al. were primarily interested in finding “non-essential” hosts, i.e. hosts that can be blocked, for example for privacy reasons, without negatively impacting the essential functionality of an IoT device. Consequently, while the total numbers are known, they only published the domain names of non-essential hosts, whereas we do not differ between essential and non-essential ones. Nevertheless, we have also included the actual domain names in Table III. The differences between the two works can be explained by the fact that our approach actively looks for hidden communication patterns, the fact that their work is focused on identifying non-essential hosts, and the reasons already mentioned above. For example, the domain name ecdinterface.philips.com does not exist anymore (in 2025), and the domain name diagnostics.meethue.com could be part of a periodic or sporadic pattern filtered out by our approach. This can also explain the empty intersections between the two works for two of the tested devices.
@@ V Related Work
Packets PingPong [11], D-interact [44] IoTa [45] De Keersmaeker [9]
Flows Mandalari [24], This work MUD [6], Hamza [10] De Keersmaeker [9], Girish [4]
@@
Among the numerous proposals for IoT device behavior models, the most widespread is probably the IETF’s MUD standard [6], as it has been standardized and backed up by multiple big players, including Cisco [46]. Subsequently, research leveraging this standard was fueled, including Hamza et al.’s solution to generate MUD profiles from network traces of IoT devices’ traffic [10]. The generated profiles, however, suffer from the shortcomings inherent to the MUD standard, namely, among others, the lack of support for protocols other than IP (v4 & v6) on layer 3, and TCP, UDP and ICMP on layer 4. De Keersmaeker et al. [9] proposed a syntax to express the intended network behavior of devices, inspired by MUD, while overcoming some of its shortcomings, and including support for traffic related to cross-device interactions.
@@ VI Conclusion
As future work, we would like to apply our methodology to scenarios where multiple Smart Home devices are involved in a single interaction. This is conceptually supported by our framework, but will require an extension of the mechanisms we use for event triggering. In this case, approaches to further automate the latter should also be investigated.
-A Statement on ethical issues
No ethical issues were raised by this work. Our device profiling activities are in line with the EU legislation on reverse engineering [47, 48]. All experiments were performed in a controlled environment, and did not involve any user personal data or other privacy concerns.
-B Additional results
Fig. 11 shows the count of unique Flow IDs discovered by our framework for each investigated event, split into first-level and hidden Flow IDs. The x-axis pertains to the event ID, as provided in Table I.
Figure 11: Number of unique Flow IDs per event.

### v5

@@ I Introduction
In this paper, we present a framework, mostly automated and requiring little prior configuration, to comprehensively uncover the communication patterns of a Smart Home device. For each type of interaction with the device, it first observes the traffic flows that occur in an unconstrained network and builds a set of flow descriptors from them. Such descriptors contain Internet and transport-layer information, as well as selected, available application-layer data, such as domain names. The system then iteratively blocks flows corresponding to certain descriptors and repeats the interaction with the intent of making new flows appear. The algorithm selecting which flows to block and the sequence of blocking experiments is not straightforward. Our aim is to discover as many hidden flows as possible but keep the number of experiments under control. To do this efficiently and not end up in an infinite loop where the same patterns appear over and over again, we store the found descriptors in a tree that we traverse using a breadth-first strategy, and prune using an ad hoc heuristic. The result is a set of multi-level, tree-shaped profiles (one for each type of interaction) for the device that describe the device’s default and alternative communication patterns that we then explore to assess the robustness of the device’s network communications.
@@
  • We develop and implement a framework, modular and easily extensible, which instruments Smart Home devices and extracts multi-level profiles for them, in a mostly automated way. The core of the framework is a new algorithm that, for a given device interaction, builds an event signature, i.e., the description of flows appearing for that interaction, and then iteratively blocks previously observed flows, until the interaction fails or no new flow is discovered.
@@
  • We publish the source code of our signature extraction algorithm at https://github.com/smart-home-network-security/signature-extraction, and of our experimental framework, including the captures of the testbed’s traffic at https://github.com/smart-home-network-security/smart-home-hidden-communication-patterns.
@@ II Problem statement and scope
The terms IoT and Smart Home are quite fuzzily defined in the literature. Our work focuses on domestic devices, augmented with IoT capabilities, which exhibit a precise, goal-oriented, and limited set of functions, e.g. power plugs, light bulbs, and cameras. More powerful/general-purpose smart home devices (e.g. smart TVs or speaker hubs) do not exhibit such a precise set, and are therefore considered out of scope. Speakers usually act as an intermediary between the user and the end device, enabling the user with voice control. Additionally, such devices allow installing third-party apps which provide them with limitless capabilities. This makes the network behavior of such devices very complex and customizable, rendering an automated analysis with our framework practically unfeasible. Indeed, as the set of functions is unlimited and not known a priori, a single device can showcase an unbounded amount of different tree profiles, as one profile pertains to one function.
@@ III-A General workflow
Fig. 2 gives an overview on the workflow of our methodology to obtain a profile for a Smart Home device under a certain type of interaction. An interaction can require the usage of a triggering source, such as a companion app installed on a smartphone. In practice, a Smart Home device may support different interactions (e.g., for a smart lamp: switching on/off, changing its color, etc.) and triggering sources, and the workflow described here must be repeated for each of them. The workflow consists of multiple steps summarized below. Besides initial manual configuration in Step 1, the rest of the workflow is automated.
@@ III-B Step 1: Traffic capturing
Similarly to PingPong [11] and IoTAthena [15], we start capturing the network traffic, perform the interaction, then stop the traffic capture after a predefined duration $d$. We repeat each interaction and the traffic capturing $m$ times. Our framework automates interactions triggered by a companion app by leveraging the Android Debug Bridge (ADB) tool [22] to generate touch events on the smartphone running the companion app. The only manual configuration required by the framework is to plug in the device under test and to connect it to the LAN, as well as setting up the framework with the smartphone’s screen coordinates to issue the touch events. We also consider power-cycling the Smart home device as an interaction (called boot interaction in the following). To automate the boot interaction, we plug the device in a smart power outlet controlled by our framework.
@@ III-D2 Tree traversal and pruning
Secondly, we prune branches of the tree which will likely not provide new information, i.e. no new Flow IDs, by applying an ad hoc pruning heuristic. To determine which heuristic we can apply without losing any information, we ran preliminary experiments with one of our testbed’s devices, the TP-Link HS110 smart plug [27]. We observed that, for any two nodes with the same Flow ID, their children were always identical. We conclude it is likely that no new information will be found by processing a node identical to one which has already been processed, regardless of its position in the tree. Therefore, we adopt the node pruning heuristic, illustrated in Fig. 3: a node is pruned if an equivalent node, i.e., a node with the same Flow ID, has already been processed. Additional insight concerning this preliminary experiment is given in appendix -C.
@@ III-D3 How to block packets matching a Flow ID
As explained, a Flow ID comprises traffic features from multiple networking layers, such as hostnames, port numbers, and application layer information. The presence of application layer information in the Flow ID requires a firewall that is able to match and reject traffic based on information found at that layer. A simple firewall, such as the default Linux kernel firewall NFTables [28] (successor of the well-known IPTables) is not sufficient. Instead, we leverage the open-source Smart Home firewall we developed as previous work [9]. It is based on NFTables, while enhancing its capabilities to match additional application-layer protocols. However, this firewall can only work with allow rules, while our profiling algorithm is based on blocking rules. Therefore, we modified the firewall, turning it into a deny-list firewall. In our prototype implementation of the profiling algorithm, we translate the list of Flow IDs into firewall blocking rules and let the firewall enforce them for the next batch of experiments. The code of the modified firewall is available at https://github.com/smart-home-network-security/firewall-blocklist.
@@ IV-A Experiment setup
In a real Smart Home network, a user could choose among various apps to control their devices, including the device’s official companion app (from its manufacturer or vendor), a third-party app, or a Smart Home automation platform which provides unified control over devices from different manufacturers. Such different home automation methods can trigger diverging, but valid, communication patterns. To cover such cases, for all interactions other than booting, we derive up to three usage scenarios:
@@
Woox R5024 (Tuya)
[35]
/ boot 4
@@
Xiaomi MJSXJ02CM
[36]
/ boot 5
@@
Philips Hue White and
Color lamp A60 $\dagger$}
[39]
/ boot 8
@@
Alecto Smart-Bulb10
(Tuya) [40]
/ boot 9
@@
TABLE I: Devices in our testbed, with corresponding apps and interactions. $\dagger$} indicates Zigbee devices, for which the hub traffic was considered.
The complete device references are listed in Table I, along with the corresponding smartphone apps and interactions instrumented. The numerical identifier used as x-axis for some subsequent plots pertains to the Event ID listed in the table, which is a unique number assigned by us to each of the tested combinations of device, app, and interaction. As explained, we only consider IP traffic, whether over Ethernet or Wi-Fi. Therefore, for the two devices communicating using Zigbee, we profile the traffic issued by their hub, i.e. the Hue bridge for the Hue lamp, and the SmartThings hub for the SmartThings Outlet, respectively. For the boot interaction experiments over those devices, we synchronously power-cycle both the device itself and its hub.
@@ IV-B Example signature tree
Figure 5: Event signature tree for the TP-Link HS110’s toggle event. $\leftrightarrow$ (resp. $\rightarrow$) denotes bidirectional (resp. unidirectional) flows. When indicated, the port is associated with its host.
@@
    between the phone and the plug’s TCP port 9999;
@@
    between the phone and the plug’s UDP port 9999;
@@
    from the phone to the broadcast address’ (255.255.255.255) UDP port 9999;
@@ IV-C Assessing device event robustness
The experiments in our testbed of devices result in the extraction of 254 unique Flow IDs, of which 70 are hidden (i.e. 27.56%). Their distribution over the 36 tested event configurations is shown in Fig. 11 in the appendix. Hidden Flow IDs embody the alternative communication strategies that a device might use if the default one fails, effectively enhancing the device’s robustness to network instability (remember that we only keep Flow IDs associated with successful event executions). We define the robustness score as the count of hidden Flow IDs. Recall that a hidden Flow ID is one which only appears in the tree deeper than the first level, i.e. resulting from blocking specific flows. We compute this metric for each tested event; Fig. 6 shows the results.
Figure 6: Robustness score. The x-axis shows the Event ID.
@@
Figure 7: Robustness score for all events, grouped per device category, app, and manufacturer. Each marker represents the robustness score of one event. ST: SmartThings.
@@ IV-D A closer look at DNS data
0.25Refer to caption Refer to caption OcFFL
Figure 8: Domain names / DNS resolvers contacted for each boot event. The x-axis shows the Event ID.
@@ IV-D2 DNS resolvers
Out of the 17 DNS resolvers contacted, only two are hidden resolvers. One of them is simply the LAN gateway (192.168.1.1), meaning the only real new DNS resolver discovered at a level deeper than one is the Chinese public DNS resolver 114.114.114.114, contacted by the Xiaomi camera when the LAN gateway is not responsive. This result shows that, regarding domain name resolving, Smart Home devices still lack robustness. Indeed, every device should try contacting a backup resolver if the primary one fails, as this has been made easy thanks to publicly available resolvers such as Cloudflare’s 1.1.1.1 or Google’s 8.8.8.8.
@@ IV-E Analysis of the node pruning heuristic
Figure 9: Pruned Flow IDs. The x-axis shows the Event ID.
@@ IV-F Comparison with related work
TP-Link HS110
boot & toggle
5 4
use1-api.tplinkra.com
euw1-api.tplinkra.com
euw1-device-telemetry-gw.iot.i.tplinknbu
n-wap.tplinkcloud.com
n-devs.tplinkcloud.com
n-deventry.tplinkcloud.com
SmartThings Outlet
boot
4 3 /
hub.smartthings.com
usher.connect.smartthings.com
connectivity.smartthings.com
dc2-eu01-euwest1.connect.smartthings.com
fw-update2.smartthings.com
Philips Hue lamp
boot
5 4 /
mqtt-eu-01.iot.meethue.com
time.meethue.com
ntp2.aliyun.com
ntp4.aliyun.com
time4.google.com
diagnostics.meethue.com
ecdinterface.philips.com
@@
A direct comparison of the domain names beyond the mere numbers is unfortunately not possible. This is because Mandalari et al. were primarily interested in finding “non-essential” hosts, i.e. hosts that can be blocked, for example for privacy reasons, without negatively impacting the essential functionality of an IoT device. Consequently, while the total numbers are known, they only published the domain names of non-essential hosts, whereas we do not differ between essential and non-essential ones. Nevertheless, we have also included the actual domain names in Table III. The differences between the two works can be explained by the fact that our approach actively looks for hidden communication patterns, the fact that their work is focused on identifying non-essential hosts, and the reasons already mentioned above. For example, the domain name ecdinterface.philips.com does not exist anymore (in 2025), while diagnostics.meethue.com could be part of a periodic or sporadic pattern filtered out by our approach. This can also explain the empty intersections between the two works for two of the tested devices.
@@ V Related Work
Packets
PingPong [11],
D-interact [44]
IoTa [45] Previous work [9]
Flows
Mandalari [24],
This work
MUD [6],
Hamza [10]
Previous work [9],
Girish [4]
@@
Among the numerous proposals for IoT device behavior models, the most widespread is probably the IETF’s MUD standard [6], as it has been standardized and backed up by multiple big players, including Cisco [46]. Subsequently, research leveraging this standard was fueled, including Hamza et al.’s solution to generate MUD profiles from network traces of IoT devices’ traffic [10]. The generated profiles, however, suffer from the shortcomings inherent to the MUD standard, namely, among others, the lack of support for protocols other than IP (v4 & v6) on layer 3, and TCP, UDP and ICMP on layer 4. In previous work [9] proposed a syntax to express the intended network behavior of devices, inspired by MUD, while overcoming some of its shortcomings, and including support for traffic related to cross-device interactions.
@@
Finally, the concept of chaos engineering [47] is related to our approach. It consists in injecting faults (e.g. network loss or latency) or heavy load (e.g. on CPU, memory, etc.). While our approach shares some concepts with chaos engineering, its novelty resides in the fact that we are the first to apply it to this extent in the context of smart homes. Moreover, today, no chaos engineering tool provides control over network traffic at a protocol level as precise as us, by blocking traffic based on specific fields. The most precise is Steadybit [48], which allows blocking TCP altogether.
@@ VI Conclusion
As future work, we would like to apply our methodology to scenarios where multiple Smart Home devices are involved in a single interaction. This is conceptually supported by our framework, but will require an extension of the mechanisms we use for event triggering. In this case, approaches to further automate the latter should also be investigated. Besides, our robustness analysis only focused on binary traffic verdict, i.e. either the traffic passes or is blocked. To further extend our robustness analysis, we could consider other traffic disturbances, such as limiting the traffic rate and inserting traffic bursts. A study covering such bursts would be transversal to ours, focusing more on robustness with respect to, e.g., DoS attacks.
Acknowledgments
F. De Keersmaeker is an F.R.S.-FNRS Research Fellow. R. Sadre is supported by the Walloon Belgian region project CyberExcellence. C. Pelsser is supported by the Cisco University Research Program Fund, an advised fund of Silicon Valley Foundation, and the Win4Collective CARAPACE project. All icons are from flaticon.com.